AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 577     Entry time: Sun Jun 2 22:46:47 2024
Author: JM, EM, NP 
Subject: Run24 

23:47

Run 24 started - Beam on target

Beam energy is 1.5 MeV.

Horizontal beam position = 7.0 mm, vertical beam position = -1.0 mm.

Target density ~ 1e11 at/cm2 (Still Fluctuating, after temperature increase)

Beam intensity at inj ~ 2.5e7

Detectors are positioned IN. Left arm = 62.71, right arm =  64.60 mm

Histograms and stats zero'd, monitoring code restarted.

Beam in the ring ~ 10 minutes.

Pressure in YR09 E target ~ 5.2e-11 mbar.

 

Merger OK, Tape server OK

ADC data item statistics looks as before. OK

FEE temperatures look OK

System checks OK

Leakage current stable

 

00:23

 

Target density ~4.65e11 atoms/cm2

Beam intensity at inj ~ 2.7e7

Beam intensity dropped to 2.1e7 in 10 minutes. Proportionall this is much greater than before. 

Run 24 stopped.

 

Attachment 1: Screenshot_from_2024-06-02_23-53-00.png  116 kB  Uploaded Sun Jun 2 23:55:39 2024  | Hide | Hide all
Screenshot_from_2024-06-02_23-53-00.png
Attachment 2: Screenshot_from_2024-06-02_23-53-47.png  164 kB  Uploaded Sun Jun 2 23:55:46 2024  | Hide | Hide all
Screenshot_from_2024-06-02_23-53-47.png
Attachment 3: Screenshot_from_2024-06-02_23-56-22.png  426 kB  Uploaded Sun Jun 2 23:56:27 2024  | Hide | Hide all
Screenshot_from_2024-06-02_23-56-22.png
ELOG V3.1.4-unknown